If a bus travels 18 miles in 30 minutes, how far will it travel in 10 Minutes?
Answer by unlockmath(1688)

Hello,
We can set up a proportion as follows:
18/30=x/10
Multiply by 10 both sides:
180/30=x
x=6 miles
Make sense?
RJ
www.math-unlock.com